# Sora: A Review on Background, Technology, Limitations, and Opportunities of Large Vision Models

### v2

@@ Abstract
Warning: This is not an official technical report from OpenAI.
@@ 3.3.1 Diffusion Transformer
Like Masked AutoEncoder (MAE) [33], Masked Diffusion Transformer (MDT) [55] incorporates mask latent modeling into the diffusion process to explicitly enhance contextual relation learning among object semantic parts in image synthesis. Specifically, as shown in Figure 12, MDT uses a side-interpolated for an additional masked token reconstruction task during training to boost the training efficiency and learn powerful context-aware positional embedding for inference. Compared to DiT [4], MDT achieves better performance and faster learning speed. Instead of using AdaLN (i.e., shifting and scaling) for time-conditioning modeling, Hatamizadeh et al. [56] introduce Diffusion Vision Transformers (DiffiT), which uses a time-dependent self-attention (TMSA) module to model dynamic denoising behavior over sampling time steps. Besides, DiffiT uses two hybrid hierarchical architectures for efficient denoising in the pixel space and the latent space, respectively, and achieves new state-of-the-art results across various generation tasks. Overall, these studies show promising results in employing vision transformers for image latent diffusion, paving the way for future studies for other modalities.
@@
Video Diffusion Transformer. Building upon the foundational works in text-to-image (T2I) diffusion models, recent research has been focused on realizing the potential of diffusion transformers for text-to-video (T2V) generation tasks. Due to the temporal nature of videos, key challenges for applying DiTs in the video domain are: i) how to compress the video spatially and temporally to a latent space for efficient denoising; ii) how to convert the compressed latent to patches and feed them to the transformer; and iii) how to handle long-range temporal and spatial dependencies and ensure content consistency. Please refer to Section 3.2.3 for the first challenge. In this Section, we focus our discussion on transformer-based denoising network architectures designed to operate in the spatially and temporally compressed latent space. We give a detailed review of the two important works (Imagen Video [29] and Video LDM [36]) described in the reference list of the OpenAI Sora technique report.
Imagen Video [29], a text-to-video generation system developed by Google Research, utilizes a cascade of diffusion models, which consists of 7 sub-models that perform text-conditional video generation, spatial super-resolution, and temporal super-resolution, to transform textual prompts into high-definition videos. As shown in Figure 13, firstly, a frozen T5 text encoder generates contextual embeddings from the input text prompt. These embeddings are critical for aligning the generated video with the text prompt and are injected into all models in the cascade, in addition to the base model. Subsequently, the embedding is fed to the base model for low-resolution video generation, which is then refined by cascaded diffusion models to increase the resolution. The base video and super-resolution models use a 3D U-Net architecture in a space-time separable fashion. This architecture weaves temporal attention and convolution layers with spatial counterparts to efficiently capture inter-frame dependencies. It employs v-prediction parameterization for numerical stability and conditioning augmentation to facilitate parallel training across models. The process involves joint training on both images and videos, treating each image as a frame to leverage larger datasets, and using classifier-free guidance [57] to enhance prompt fidelity. Progressive distillation [58] is applied to streamline the sampling process, significantly reducing the computational load while maintaining perceptual quality. Combining these methods and techniques allows Imagen Video to generate videos with not only high fidelity but also remarkable controllability, as demonstrated by its ability to produce diverse videos, text animations, and content in various artistic styles.
@@
Blattmann et al. [36] propose to turn a 2D Latent Diffusion Model into a Video Latent Diffusion Model (Video LDM). They achieve this by adding some post-hoc temporal layers among the existing spatial layers into both the U-Net backbone and the VAE decoder that learns to align individual frames. These temporal layers are trained on encoded video data, while the spatial layers remain fixed, allowing the model to leverage large image datasets for pre-training. The LDM’s decoder is fine-tuned for temporal consistency in pixel space and temporally aligning diffusion model upsamplers for enhanced spatial resolution. To generate very long videos, models are trained to predict a future frame given a number of context frames, allowing for classifier-free guidance during sampling. To achieve high temporal resolution, the video synthesis process is divided into key frame generation and interpolation between these key frames. Following cascaded LDMs, a DM is used to further scale up the Video LDM outputs by 4 times, ensuring high spatial resolution while maintaining temporal consistency. This approach enables the generation of globally coherent long videos in a computationally efficient manner. Additionally, the authors demonstrate the ability to transform pre-trained image LDMs (e.g., Stable Diffusion) into text-to-video models by training only the temporal alignment layers, achieving video synthesis with resolutions up to 1280 × 2048.
@@ 3.3.2 Discussion
Cascade diffusion models for spatial and temporal up-sampling. Sora can generate high-resolution videos. By reviewing existing works and our reverse engineering, we speculate that Sora also leverages cascade diffusion model architecture [59] which is composed of a base model and many space-time refiner models. The attention modules are unlikely to be heavily used in the based diffusion model and low-resolution diffusion model, considering the high computation cost and limited performance gain of using attention machines in high-resolution cases. For spatial and temporal scene consistency, as previous works show that temporal consistency is more important than spatial consistency for video/scene generation, Sora is likely to leverage an efficient training strategy by using longer video (for temporal consistency) with lower resolution. Moreover, Sora is likely to use a $v$-parameterization diffusion model [58], considering its superior performance compared to other variants that predict the original latent $x$ or the noise $\epsilon$.
On the latent encoder. For training efficiency, most of the existing works leverage the pre-trained VAE encoder of Stable Diffusions [60, 61], a pre-trained 2D diffusion model, as an initialized model checkpoint. However, the encoder lacks the temporal compression ability. Even though some works propose to only fine-tune the decoder for handling temporal information, the decoder’s performance of dealing with video temporal data in the compressed latent space remains sub-optimal. Based on the technique report, our reverse engineering shows that, instead of using an existing pre-trained VAE encoder, it is likely that Sora uses a space-time VAE encoder, trained from scratch on video data, which performs better than existing ones with a video-orient compressed latent space.

### v3

@@ Abstract
Note: This is not an official technical report from OpenAI.
@@ 3.3.1 Diffusion Transformer
Like Masked AutoEncoder (MAE) [33], Masked Diffusion Transformer (MDT) [55] incorporates mask latent modeling into the diffusion process to explicitly enhance contextual relation learning among object semantic parts in image synthesis. Specifically, as shown in Figure 12, MDT uses a side-interpolated for an additional masked token prediction task during training to boost the training efficiency and learn powerful context-aware positional embedding for inference. Compared to DiT [4], MDT achieves better performance and faster learning speed. In MDTv2 [56], they further improve MDT with a more efficient macro network with some u-shape long-shortcuts in the encoder and dense shortcuts from encoder input in the decoder. Furthermore, they also leverage a bag of training strategies like the Adan optimizer [57], Min-SNR weighting [58], dynamic masking ratio, and an improved power-cosine weighting for classifier-free guidance [59]. MDTv2 achieves a significant boost in synthesis performance and learning speed. Instead of using AdaLN (i.e., shifting and scaling) for time-conditioning modeling, Hatamizadeh et al. [60] introduce Diffusion Vision Transformers (DiffiT), which uses a time-dependent self-attention (TMSA) module to model dynamic denoising behavior over sampling time steps. Besides, DiffiT uses two hybrid hierarchical architectures for efficient denoising in the pixel space and the latent space, respectively, and achieves new state-of-the-art results across various generation tasks. Overall, these studies show promising results in employing vision transformers for image latent diffusion, paving the way for future studies for other modalities.
@@
Video Diffusion Transformer. Building upon the foundational works in text-to-image (T2I) diffusion models, recent research has been focused on realizing the potential of diffusion transformers for text-to-video (T2V) generation tasks. Due to the temporal nature of videos, key challenges for applying DiTs in the video domain are: i) how to compress the video spatially and temporally to a latent space for efficient denoising; ii) how to convert the compressed latent to patches and feed them to the transformer; and iii) how to handle long-range temporal and spatial dependencies and ensure content consistency. Please refer to Section 3.2.3 for the first challenge. In this section, we focus our discussion on transformer-based denoising network architectures designed to operate in the spatially and temporally compressed latent space. We give a detailed review of the two important works (Imagen Video [29] and Video LDM [36]) shown in the OpenAI Sora technique report and also briefly introduce some of recent promising approaches.
Imagen Video [29], a text-to-video generation system developed by Google Research, utilizes a cascade of diffusion models, which consists of 7 sub-models that perform text-conditional video generation, spatial super-resolution, and temporal super-resolution, to transform textual prompts into high-definition videos. As shown in Figure 13, firstly, a frozen T5 text encoder generates contextual embeddings from the input text prompt. These embeddings are critical for aligning the generated video with the text prompt and are injected into all models in the cascade, in addition to the base model. Subsequently, the embedding is fed to the base model for low-resolution video generation, which is then refined by cascaded diffusion models to increase the resolution. The base video and super-resolution models use a 3D U-Net architecture in a spatial-temporal separable fashion. This architecture weaves temporal attention and convolution layers with spatial counterparts to efficiently capture inter-frame dependencies. It employs v-prediction parameterization for numerical stability and conditioning augmentation to facilitate parallel training across models. The process involves joint training on both images and videos, treating each image as a frame to leverage larger datasets, and using classifier-free guidance [59] to enhance prompt fidelity. Progressive distillation [61] is applied to streamline the sampling process, significantly reducing the computational load while maintaining perceptual quality. Combining these methods and techniques allows Imagen Video to generate videos with not only high fidelity but also remarkable controllability, as demonstrated by its ability to produce diverse videos, text animations, and content in various artistic styles.
@@
Blattmann et al. [36] propose to turn a 2D Latent Diffusion Model into a Video Latent Diffusion Model (Video LDM). They achieve this by adding some post-hoc temporal layers among the existing spatial layers into both the U-Net backbone and the VAE decoder that learns to align individual frames. These temporal layers are trained on encoded video data, while the spatial layers remain fixed, allowing the model to leverage large image datasets for pre-training. The LDM’s decoder is fine-tuned for temporal consistency in pixel space and temporally aligning diffusion model upsamplers for enhanced spatial resolution. To generate very long videos, models are trained to predict a future frame given a number of context frames, allowing for classifier-free guidance during sampling. To achieve high temporal resolution, the video synthesis process is divided into key frame generation and interpolation between these keyframes. Following cascaded LDMs, a DM is used to further scale up the Video LDM outputs by four times, ensuring high spatial resolution while maintaining temporal consistency. This approach enables the generation of globally coherent long videos in a computationally efficient manner. Additionally, the authors demonstrate the ability to transform pre-trained image LDMs (e.g., Stable Diffusion) into text-to-video models by training only the temporal alignment layers, achieving video synthesis with resolutions up to $1280\times 2048$px.
Recently, there are also some other promising works that show superior performance in high-quality and temporal-consistent video generation. Similar to Imagen Video, LAVIE [62] proposes to transform a T2I model into a T2V model with cascaded LDMs, with two key insights: i) using simply temporal self-attention and RoPE [63] can adequately capture temporal correlations inherent in video data; ii) joint image-video fine-tuning enable more stable training and avoid catastrophic forgetting. GenTron [64] adapt pre-trained class-conditioning DiTs to textual conditioning and explore different conditioning mechanisms, where they find that cross attention is a better conditioning method compared to previous adaLM-Zero for free-form textual conditioning. Furthermore, by designing motion-free masking in the inserted temporal self-attention layers, GenTron is able to do efficient image-video joint training with control on the temporal modeling ability in the video diffusion process. Window Attention Latent Transformer (W.A.L.T) [65] proposes to unify the latent space of both videos and images with 3D causal CNN and a special coding strategy for the first frame. Moreover, W.A.L.T uses non-overlapping, window-restricted spatial and spatiotemporal attention to lower the computational cost and provide flexibility for image-video joint training. Latent Diffusion Transformer (Latte) [66] conducts extensive experiments among four model variants from the perspective of decomposing spatial and temporal information in video data. Furthermore, Latte studies the effect of a wide range of training strategies and modules in a DiT-based video generation system, including conditioning mechanisms and temporal positional embedding. Instead of using separate compression for spatial and temporal dimensions in previous works, SnapVideo [67] proposes to extend EDM framework [68] for high-dimensional video data. Furthermore, SnapVideo treats images as $T$ frames videos with infinite frame rates and introduces a variable frame-rate training procedure to bridge the modality gap. For scaling up to billion-parameters size, SnapVideo uses an efficient transformer-based architecture, FITs [69], as a scalable video generator, and a model cascade consisting of a first-stage lower-solution base model ($36\times 64$px) and a high-solution upsampler ($288\times 512$px) with corrupt training strategy.
@@ 3.3.2 Discussion
One end-to-end model or cascade structure. Sora can generate high-resolution videos. By reviewing existing works and our reverse engineering, we speculate that Sora also leverages cascade diffusion model architecture [70], which is composed of a base model and many spatial-temporal refiner models. The attention modules are unlikely to be heavily used in the based diffusion model and low-resolution diffusion model, considering the high computation cost and limited performance gain of using attention machines in high-resolution cases. For spatial and temporal scene consistency, as previous works show that temporal consistency is more important than spatial consistency for video/scene generation, Sora is likely to leverage an efficient training strategy by using longer video (for temporal consistency) with lower resolution. Moreover, Sora is likely to use a $v$-parameterization diffusion model [61], considering its superior performance among others that predict the original latent $x$ or the noise $\epsilon$.
On the compression of the latent encoder. For training efficiency, most of the existing works leverage the pre-trained VAE encoder of Stable Diffusions [71, 72], a pre-trained 2D diffusion model, as an initialized model checkpoint. However, the encoder lacks the ability to compress temporally. Even though some works propose to only fine-tune the decoder for handling temporal information, the decoder’s performance of dealing with video temporal data in the compressed latent space remains sub-optimal. Based on the technique report, our reverse engineering shows that, instead of using an existing pre-trained Image VAE encoder with some post-hoc temporal tuning, it is likely that Sora uses a spatial-temporal VAE encoder, jointly trained on image and video data, to compress the spatial and temporal information altogether.